#2 The manager of a car wash must decide whether to have one or two wash lines. One line will mean a fixed cost of $6,500 a month, and two lines will mean a fixed cost of $11,000 a month. Each line would be able to process 14 cars an hour. Variable costs will be $4 per car, and revenue will be $5.95 per car. The manager projects an average demand of between 14 and 18 cars an hour. Would you recommend one or two lines? The car wash is open 300 hours a month.
Answer #1

In case of One lIne

The Expense would be = Fixed cost + Variable cost

= 6500 + 14*4*300

= 6500 + 16800

= $23300

Revenue = Number of cars processed per hour * Total hours open * Revenue per car

= 14*300*5.95

= $24990

Profit = Revenue - Expense

= (24990-23300)

= $1690

In case of two lines

Expense = 11000 + (14*2*4*300)

= 11000+33600

= $44600

Revenue = 14*2*300*5.95

= $49980

Profit = (49980-44600)

= $5380

I would recommend two lines as the profit for the same is more.